OpenStudy (anonymous):

Solve the equation: x2 + 8x + 16 = 49 Enter the smallest zero value. Use a number only.

OpenStudy (hoblos):

x2 + 8x + 16 = 49 using a^2 +2ab + b^2 = (a+b)^2 we get (x+4)^2 = 7^2 x+4 = ±7 x=3 or x=-11

OpenStudy (anonymous):

Solve the equation: (x + 1)2 = 9. Enter the largest zero value. Use a number only.

OpenStudy (hoblos):

it is just the same!! can you try it?

OpenStudy (anonymous):

2x +2 =9 2x=7 x=7/2???